A sample of 75 concrete blocks had a mean mass of 38.3 kg with a standard deviation of 0.5 kg.

An engineer claims that the mean mass is between 38.2 and 38.4 kg. With what level of confidence can this statement be made? (Express the final answer as a percent and round to two decimal places.)

The level of confidence is
Answer #1

Given data 02759,=38-3,0 20.5 confidence =38.4 Upper :. 38.3+Zx1 = 38.4 2412*0.0579=0.10 24/2=1.733 By using z-table, right t